wiki:UC20190401

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 4/01/19

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), Moire Prescott (NMSU), Mary Beth Kaiser (JHU), Chip Kobulnicky (UWy), Misty Bentz (GSU), Joanne Hughes (Seattle U), Ben Williams (UW), Eric Bellm (UW), Anne Verbiscer (UVa), Kal Kadlec (UW)


User feedback and comments from institutional representatives:

  • NMSU - nothing to report from users; Moire had a question about the ARCSAT guide camera and the Users' Manual (discussed further below)
  • Wyoming - nothing to report
  • UVa - nothing to report
  • Colorado - no report
  • Washington - their users are happy to see both sides of DIS running; thanks to observatory staff for making that possible
  • Seattle - nothing to report
  • Oklahoma - no report
  • Georgia State - a user had a half night of observing time with DIS last week and it went well; the data are of good quality. Thanks to the observatory staff for getting it working
  • NAPG - no report
  • JHU - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 3/01/19 - 3/27/19

0) Overview

Winter weather has improved with the arrival of spring, but that includes spring winds and dust. Included in this was a major wind event with some snow. This storm took out utility power on almost all of the mountain area. It took 3 days to restore power even in Cloudcroft (their highest priority). APO was without utility power for 5 days. We ran on our generator, however, on day 2 it developed a radiator leak. It took nearly all the staff to monitor and report regular status so we could be prepared in the event APO went completely dark. Having a known issue with the generator kept us closed although in fact the weather was too bad to observe most of the time.

With power issues we were very conservative on vacuum pumping DIS Blue camera so this did delay that camera coming back on-line. There were no other issues with the telescope, instruments, or computers due to these power issues.

No other significant issues

1) Telescope

The telescope had no problems.

2) Instruments

DIS: The DIS blue camera is now back in service. We did find some crossed wires in the dewar. The new getter holder makes this more likely to occur. Our current thoughts are that we will not routinely bake the getter out when we service the instrument. We will only do so if there is evidence that removal and bake-out is needed. The DIS red camera contamination is holding well. We also will not plan to regularly service the cameras until there is evidence that it is needed. We often service them during the summer regardless of need to try and avoid winter servicing.

TripleSpec is operational.

Agile is operational with very little usage. We had a minor issue with the rotator where an umbilical snagged on a bolt.

Echelle is operational. The science instrument had a intermittent shutter failure. This is a homemade shutter by Yerkes and the screws holding the solenoid were slightly loose. We were able to repair this, at least for a period of time. The new Ecam is not quite ready for prime time but the communication issue has been resolved by a high dollar USB extender.

NIC-FPS is operational. The ICC was repaired. We had issues with the terminal server being locked up when the ICC power supply died. This resulted in a fair amount of extra troubleshooting and repair time to correct.

ARCTIC is operational.

3) ARCSAT

Operational with the new guide camera system, which has been relocated on the dec axis, realigned and adjusted to improve guiding. We are still seeing issues with the old DFM electronics and commercial software and looking at alternatives for the future, since the upgrades planned for during the last engineering window could not be implemented.


Additional telescope and instrument discussion:

The weather knocked out power to the site for 5 days. Our backup generator kept power going to the site but it took a lot of effort from the staff. Therefore we are looking at options for purchasing a backup to the backup. It won’t be a cheap solution but we may need to consider it given the impact of what would happen if we were to go dark (including lost time and the fact that some of our equipment is aging and may not come back gracefully).

The bright spot in the activities over the past month was bringing DIS back online (which took longer than planned due to the power outage). The engineering staff found crossed wires inside the dewar but determined that this did not damage the chip. We think it happened during reassembly after a servicing period; this process is now more difficult due to the addition of getter material intended to reduce the contamination. Going forward we will be more conservative regarding when we service the dewar and how deeply we go into it.

The ARCES shutter problem has been repaired. It was a homemade shutter delivered from Yerkes when we received ARCES, so we are not confident that the problem will not recur. Progress was made on the new Ecam; it is coming along. The ICC for NICFPS is now working.

Regarding ARCSAT, we did not make as much progress during engineering as we would have liked. Bill put together a list of options that the telescope management team is talking about; these range from leaving things as they are to small upgrades to a substantial investment. Nothing will be decided until after the Strategic Planning meeting when this can be discussed in more detail. To address Moire’s question, Bill replied that the ARCSAT User's Manual does not currently discuss any procedures for guiding. He will update it within the next several weeks.


Q1 ARCSAT scheduling:

The last week in the Q1 schedule is now taken.


Q2 3.5-m scheduling:

The schedule is now out, although there will be a few minor adjustments coming. We are already getting requests for OPEN time -- even for late in the quarter -- so if people are planning to ask for some OPEN time please let us know soon.


Q2 ARCSAT scheduling:

We received the Q2 ARCSAT proposals last week, and the schedule will be posted soon.


NA2 Baffle Project:

John Wilson had to go back and make some modifications to the proposed baffle design now that we opened the FOV a bit (see discussion in Users Committee minutes from February 4, 2019). This caused a slight delay in sending an RFP to vendors for the stray light analysis. The baffle requires a slot in it to accommodate the NA2 guider FOV and John has been examining the pros and cons of opening it up more (impact on science instruments vs. the guider). Work on this project will proceed independent of the summer shutdown. We will be requesting users' assistance in helping to test and characterize the new FOV, and we will be discussing whether there are any current baffle data that we want to get but don’t already have. These data will be important to use as a baseline for comparison with the data that will be acquired after the new baffle is installed. We will revisit this topic next month.


KOSMOS Update:

Kal reported on behalf of the UW team that most things are on schedule. The designs are already done and now they are sending things out to get machined. They have a group working on silicon slits for the slit viewer. The dewar is moving along. Lots of testing will be required once they start assembling; that will determine whether or not they can achieve the planned schedule for June commissioning.


Response to Call for White Papers for 2020 Decadal Survey

We did not end up submitting a science white paper, but Nancy did submit a Notice of Intent for the second deadline for White Papers on Activities, Projects, or State of the Profession Considerations. The deadline for these white papers has not yet been announced (NAS web site indicates a tentative deadline of July 1). Nancy will continue to move this forward and reach back out to other observatory directors to see whether others wish to participate in this as well.


Discussion of 25th Anniversary / Strategic Planning meeting in 2019

Nancy provided some key information concerning the upcoming Strategic Planning Meeting in June:

  • Dates: June 11-12, 2019
  • Location: Sunspot Astronomy and Visitors Center
  • Lodging: Rooms will be blocked off at The Lodge in Cloudcroft
  • Attendance: is strongly requested/encouraged from all stakeholders (at least one from each institution!)
  • Travel subsidy: Nancy will cover all meals and lodging using her Director's Discretionary Funds. Any remaining DD funds will be used to help defray the travel costs for attendees.
  • Agenda: coming soon.
  • Short Term Request: discuss this with your department members and identify who will be able to attend

ACTION ITEMS:

Open action items from previous meetings:

  • Misty Bentz and John Bally will upload their NICFPS data reduction scripts to the wiki. STATUS:OPEN. Misty submitted her scripts to Nancy; she will upload them to the wiki. John still needs to submit his.
  • Ben Williams will ask Anjum whether she had any Agile data reduction documentation she could upload to the wiki. Ben got Anjum's new email address and contacted her about this. Although she has been pretty inactive lately, she said that she is interested in supplying something but she hasn’t sent anything yet. Ben will send her a reminder. STATUS: OPEN. Ben emailed her again but has not received a response. Nancy will also contact one of her former graduate students who used Agile extensively to see if he can supply any scripts.
  • Nancy will disseminate the TEWG final report. STATUS: OPEN.
  • Nancy will distribute redacted slides from her ARC BoG presentation. STATUS: OPEN. There was a discussion of where and how to them. The group agreed that they should be emailed to 35m-general and posted on the wiki.
  • Nancy will coordinate Astro2020 science white paper submission based on telecon tomorrow. STATUS: CLOSED - see above.
  • Nancy will disseminate information concerning the date/location of the Strategic Planning meeting in June 2019. STATUS: OPEN

New action items from this meeting:

  • Bill will update the ARCSAT User's Manual to include instructions for guiding.
  • All UC reps should work with their department members to identify who will attend the Strategic Planning Meeting in June.

All Other Business:

We discussed the upcoming summer shutdown. It will be very short this year, starting on July 15 and returning to science on or about July 25. This means there will be one week when the telescope is completely down and serviced, and up to another week of returning to science, including on-sky testing. The planned activities for this summer shutdown are fairly small, routine tasks: cleaning and washing all 3 mirrors, servicing the encoders, with as minimal breakdown of the telescope as possible. There will be very little instrument servicing this time. Recall that last year’s shutdown was longer than normal because we had extra coatings put on M2 and M3. Those coatings are holding up well, so aside from washing the mirrors we don’t need to do more this year.


Next meeting:

The next Users Committee telecon will be on Tuesday May 7, 2019, at 10:00 am MDT.

Note: See TracWiki for help on using the wiki.